OpenStudy (anonymous):

Find the slope and y-intercept of the line. y= 6/7x - 7

OpenStudy (godlovesme):

y=mx+b m-slope b-y-intercept make sense?

Nnesha (nnesha):

like i said before slope intercept form equation is y = mx + B so what is slope

OpenStudy (anonymous):

So remember to always get it into y=mx+b :P

Nnesha (nnesha):

@Bigtim10 so now this time tell me what is your slope suppose to be :) :) :) at least try :) :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Is the slope 6/7

OpenStudy (anonymous):

yes

Nnesha (nnesha):

yep

OpenStudy (godlovesme):

yes :) :))))

Nnesha (nnesha):

what is y-intercept ???

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

y intercept is where y intercepts with another y

OpenStudy (anonymous):

(0,-7)

OpenStudy (anonymous):

Nice job. Which one in the ordered pair though?

OpenStudy (anonymous):

-7

Nnesha (nnesha):

there you goo \(\color{blue}{GOOD :) JOB}\) i knew it you gonna do it

OpenStudy (anonymous):

Thanks for the help

Nnesha (nnesha):

\(\color{green}{MY :) PLEASURE :)}\)

OpenStudy (godlovesme):

great job!!! 

OpenStudy (godlovesme):

anytime ^.^ and dat face doe :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Nice job, @Bigtim10 :)

OpenStudy (anonymous):

Thanks